# A Survey on Compositional Learning of AI Models: Theoretical and Experimental Practices

### v1

@@ Abstract
Compositional learning, mastering the ability to combine basic concepts and construct more intricate ones, is crucial for human cognition, especially in human language comprehension and visual perception. This notion is tightly connected to generalization over unobserved situations. Despite its integral role in intelligence, there is a lack of systematic theoretical and experimental research methodologies, making it difficult to analyze the compositional learning abilities of computational models. In this paper, we survey the literature on compositional learning of AI models and the connections made to cognitive studies. We identify abstract concepts of compositionality in cognitive and linguistic studies and connect these to the computational challenges faced by language and vision models in compositional reasoning. We overview the formal definitions, tasks, evaluation benchmarks, variety of computational models, and theoretical findings. We cover modern studies on large language models to provide a deeper understanding of the cutting-edge compositional capabilities exhibited by state-of-the-art AI models and pinpoint important directions for future research.
@@ 1 Introduction
The compositional learning and reasoning of an intelligent agent refers to the ability to understand and manipulate complex structures by decomposing them into simpler parts and composing parts to form new complex concepts with a coherent understanding. This ability is a key factor in generalizing learning to unobserved situations [40]. Compositional learning in intelligent systems is cognitively motivated since humans learn compositionally [59]. Researchers have examined this phenomenon from cognitive, linguistic, and psychological perspectives [102, 20].
@@
The formal notion of compositionality originated from natural language and semantics, with various theories and arguments that elaborate on this concept. The principle of compositionality [85, 44, 77] is defined as "The meaning of a whole is a function of the meanings of the parts and of the way they are syntactically combined" with three general methods- new meanings, new basic parts, and new constructions. One of the earliest formalizations of compositionality was grounded in grammar trees when cognitive scientists proposed an "information processing" approach to create a model of the mind [114]. The birth of modern cognitive science happened following the proposal of phrase structure and transformational grammar [10]. Compositional understanding of linguistic constructs has multiple aspects [72, 63]. For example, a nesting description such as "The black tall woman on the left of the car" conveys the intersection of multiple adjectives and spatial relations. Thus, the composition is defined as the intersection of multiple concepts. However, there are cases in which the direct intersection is not applicable, and the meaning should be inferred from the concepts in the global context, such as recognizing the sentiment of the following sentence "The pizza is so good, I hate this place!". Despite natural language being a prominent manifestation of compositionality, this can be expanded to other areas of human intelligence such as vision [98]. The same notion of intersection, as well as part-whole compositions, is essential for visual intelligence.
@@
Compositional learning is important in complex tasks where high-level goals must be decomposed into smaller subgoals and plans, for example, when instructing an agent to navigate from one point to another [100]. From the computational modeling perspective, traditionally formal grammars have been the means to address the compositional understanding of various modalities primarily in language and extended to vision [26, 38]. While these models inherently address compositional structures, using them alone, that is, parsing raw and noisy data into a structure with manually designed grammars will be brittle in real-world situations. Our study focuses on data-driven approaches based on artificial neural networks and the combined paradigms of neuro-symbolic techniques. Several studies provide both experimental and theoretical analyses, indicating the competitiveness of these models in expressing compositional structures such as context-free grammars [104, 103]. Their expressive power added to the robustness in dealing with noisy data makes the neural techniques applicable to realistic situations.
@@
In this paper, we examine multiple aspects of compositional learning, including cognitive aspects, computational models, and evaluation paradigms in both theory and practice. Figure 1 shows the scope and structure of our survey, covering three main topics. Firstly, we overview the different cognitive aspects that define abstract tasks for compositionality and connect them to the existing datasets in the AI community. Secondly, we overview two evaluation approaches- theoretical and experimental. The theoretical analyses examine various computational models, investigating their expressivity, capacity for compositional learning, and analyzing the generalizations to unobserved situations. Experimental analyses include benchmarks set by datasets, created primarily to highlight the core challenges of compositional learning for language and vision understanding. Lastly, we overview the models that aim to address compositional learning, divided into categories of basic neural architectures, large language models, and customized architectures including neuro-symbolic models.
Figure 1: Outline of covered concepts in this survey, color coded to suggest research coverage and work required
@@ 2 Compositional Learning Facets
Compositionality is one important aspect of generalization as a whole [40]. Cognitive Science and Linguistics literature have identified broad categorizations of tasks that define compositionality and can be used to evaluate the compositional reasoning of models. The foundations of human natural language lie in compositionality. A commonly used task categorization, derived from reformulated theoretically grounded tests from [39], defines five main metrics of compositionality: systematicity, productivity, substitutivity, localism, and overgeneralization [14].
@@ 2.1 Systematicity or Novel Composition
Systematicity is one of the most commonly used notions of compositionality in evaluating the performance of architectures. It has been defined as the ability to systematically recombine known parts and rules [14]. It derives directly from the commonly accepted definition of composition, which is the formation of compound expressions as a function of simpler ones [85]. Systematicity is a standard concept in cognitive science research on building cognitive architectures that tried to formalize the human thinking process [19]. The ability to syntactically combine known elements to form new or "unseen" expressions is an integral test when evaluating a model’s ability to reason compositionally. It is also called Distribution Based Compositionality Assessment [49], where two principles are defined: one is to make sure the distribution of atoms is similar in both training and test sets, while another is to ensure the distribution of compounds is different.
2.2 Productivity or Length Generalization
Another commonly explored test for compositionality is length generalization or productivity, as defined in [39]. In this evaluation, models are tested on their performance with expressions or sequences that are longer than training data. Longer input sequences may be recursive or nested versions of seen phrases in the case of natural language inputs [51].
2.3 Substitutivity or Synonymity
Substitutivity is another form of evaluation defined in [14], which evaluates model performance on the introduction of synonyms in expressions. For example, testing a model on the translation of the same sentence, switching between synonymous words such as donut/doughnut or aubergine/eggplant. This is one of the lesser-tested axioms of compositionality.
2.4 Localism
Another nuance of compositionality is the notion of global versus local composition. According to the principle of compositionality, the locality of the composition operator can vary. The meaning of a complex expression can depend solely on the meaning of its immediate parts (local composition) or the global structure of the context. Localism can be tested by analyzing the meaning a model assigns to a standalone compound versus when that compound is part of a larger expression.
2.5 Overgeneralization
Overgeneralization, as defined in [14], evaluates how much a model prefers an exception versus a rule. The term is originally used in language acquisition literature, also known as overregularization. A well-known example of this is the past-tense debate [71] in language. This property can be evaluated by testing a model on exceptions of a usual rule in the training data and seeing if the model has over-fitted the training samples [39]. Another example of this task is translating idioms where the meaning of sentences is "exceptions" to usual rules. In this scenario, a model can get better performance by considering those sentences in a global sense that is looking at the bigger picture, such as context from placement in a compound, instead of trying to evaluate the meaning locally, such as by isolating the phrase.
2.6 Compositionality as Function Properties
In [96], compositional functions are defined with several components with a computation directed acyclic graph (cDAG) at the core. Their formal definitions facilitate the evaluation of compositional properties of the learning models (i.e. compositional functions). They relate their defined structure to the learning models’ expressivity and sample complexity. In this system, Systematicity can be thought of as the expressivity of a compositional function as low entropy program (for example, decision tree) versus a high entropy program (for example, transformer). Productivity, in simple terms, can be interpreted as whether a compositional function is recursive. Substitutivity tests whether a compositional function respects important abstractions and can be factored over them. Localism measures the stability of a compositional function against local changes, where the structure of the function’s elements affects the importance of the level of locality respected. Overgeneralization is the extent of compression of a function, where a function might have a general rule but have exceptions to those in special cases. [^1] Thanks to Tim Klinger for providing these insightful interpretations.
3 Abstract Tasks and Datasets
Here, we categorize the datasets based on their type of compositionality mentioned in Section 2. Table 1 points to a list of important datasets we have surveyed. We describe some of them below.
3.1 Systematicity or Novel Composition
@@
CREPE. This is a Compositional REPresentation Evaluation benchmark (CREPE) [69]. The dataset is synthesized and includes multiple splits, one of which relates to Systematicity. The main task setting is that, given an image, the model needs to identify an appropriate text caption describing it among multiple given choices. This systematicity challenge tests whether the model can systematically generate new combinations of seen atomic concepts during training. For example, "Crepe on a skillet" is never observed in the training while Crepe and skillet are observed separately in different contexts.
@@
SCAN. The task is to navigate in a two-dimensional grid world based on natural language instructions. It is the Simplified version of the CommAI Navigation tasks (SCAN) [58, 74]. One of the proposed experiments in SCAN evaluates the model’s compositional generalization across primitive commands. Specific compounds are excluded from training where the model has seen the primitives and similar compound structures. Then, these unseen compounds are tested during testing.
@@
gSCAN. The task is to navigate in a two-dimensional grid world, based on natural language instructions [97], which is a grounded version of SCAN (gSCAN). It includes 9 splits A-I. Categories B to H present tasks that form tests for systematicity (B,C- Novel Composition of Object Properties, D- Novel Direction, E- Novel Contextual References, F- Novel composition of actions and arguments, G,H- Novel Adverbs). Each split focuses on some form of novel composition of known concepts.
@@
PCFG SET. PCFG SET stands for Probabilistic Context Free Grammar String Edit Task. It is an artificial translation task where sequences produced by probabilistic context-free grammar need to be translated into sequences representing their meaning [39]. The output sequences can be constructed recursively using specified string edit operations applied to the input sequence e.g.the input ‘repeat ABC‘ will be mapped to the sequence ‘ABCABC‘. The systematicity test uses a combination of words $a$ and $b$ in the input where the model has seen $a$ but not $b$ and vice versa. However, the combination of $a$ and $b$ is plausible in the corpus.
@@
COGS. The data is designed for the so-called Compositional Generalization Challenge (COGS) [51]. The task is a kind of semantic parsing based on a fragment of English where the models need to determine a formal meaning representation of the input English sentences. There are 4 categories in COGS that test some form of systematicity including Novel Combination of Familiar Primitives and Grammatical Roles, Novel Combination of Modified Phrases and Grammatical Roles, Verb Argument Structure Alteration, and Verb Class.

### v2

@@ Abstract
Compositional learning, mastering the ability to combine basic concepts and construct more intricate ones, is crucial for human cognition, especially in human language comprehension and visual perception. This notion is tightly connected to generalization over unobserved situations. Despite its integral role in intelligence, there is a lack of systematic theoretical and experimental research methodologies, making it difficult to analyze the compositional learning abilities of computational models. In this paper, we survey the literature on compositional learning of AI models and the connections made to cognitive studies. We identify abstract concepts of compositionality in cognitive and linguistic studies and connect these to the computational challenges faced by language and vision models in compositional reasoning. We overview the formal definitions, tasks, evaluation benchmarks, various computational models, and theoretical findings. Our primary focus is on linguistic benchmarks and combining language and vision, though there is a large amount of research on compositional concept learning in the computer vision community alone. We cover modern studies on large language models to provide a deeper understanding of the cutting-edge compositional capabilities exhibited by state-of-the-art AI models and pinpoint important directions for future research.
@@ 1 Introduction
The compositional learning and reasoning of an intelligent agent refers to the ability to understand and manipulate complex structures by decomposing them into simpler parts and composing parts to form new complex concepts with a coherent understanding. This ability is a key factor in generalizing learning to unobserved situations (Hupkes et al., 2023). Compositional learning in intelligent systems is cognitively motivated since humans learn compositionally (Lake et al., 2019). Researchers have examined this phenomenon from cognitive, linguistic, and psychological perspectives (Shepard, 1987; Frankland & Greene, 2020).
@@
The formal notion of compositionality originated from natural language and semantics, with various theories and arguments that elaborate on this concept. The principle of compositionality (Partee, 2004; Janssen & Partee, 1997; Montague, 1974) is defined as “The meaning of a whole is a function of the meanings of the parts and of the way they are syntactically combined” with three general methods- new meanings, new basic parts, and new constructions. One of the earliest formalizations of compositionality was grounded in grammar trees when cognitive scientists proposed an information processing approach to create a model of the mind (Thagard, 2023). The birth of modern cognitive science happened following the proposal of phrase structure and transformational grammar (Chomsky, 1965). Compositional understanding of linguistic constructs has multiple aspects (Marelli et al., 2014; Li et al., 2021). For example, a nesting description such as “The black tall woman on the left of the car” conveys the intersection of multiple adjectives and spatial relations. Thus, the composition is defined as the intersection of multiple concepts. However, there are cases in which the direct intersection is not applicable, and the meaning should be inferred from the concepts in the global context, such as recognizing the sentiment of the following sentence “The pizza is so good, I hate this place!” Despite the natural language being a prominent manifestation of compositionality, this can be expanded to other areas of human intelligence, such as vision (Saffran et al., 2007). The same notion of intersection, as well as part-whole compositions, is essential for visual intelligence.
@@
Compositional learning is important in complex tasks where high-level goals must be decomposed into smaller subgoals and plans, for example, when instructing an agent to navigate from one point to another (Schmidhuber, 1990). From the computational modeling perspective, traditionally formal grammars have been the means to address the compositional understanding of various modalities primarily in language and extended to vision (Girshick et al., 2011; Hong et al., 2021). While symbolic models inherently address compositional structures, using them alone, that is, parsing raw and noisy data into a structure with manually designed grammars will be brittle in real-world situations. Our study focuses on data-driven approaches based on artificial neural networks and the combined paradigms of neurosymbolic techniques (e.g. see Rajaby Faghihi et al. (2021); Premsri & Kordjamshidi (2024)). Several studies provide both experimental and theoretical analyses, indicating the competitiveness of the neural models in expressing compositional structures such as context-free grammars (Siegelmann & Sontag, 1995; Shi et al., 2021). Their expressive power added to the robustness in dealing with noisy data makes the neural techniques applicable to realistic situations.
@@
In this paper, we examine multiple aspects of compositional learning, including compositional learning facets, datasets, computational models, and evaluation paradigms in both theory and practice. Figure 1 shows the scope and structure of our survey, covering four main topics, that is, Compositional Learning Facets, Datasets, Compositional Learning Models, and Evaluation. For Compositional Learning Facets, we overview the different measures of compositionality rooted in cognitive science that define abstract tasks for compositionality and connect them to the other adjacent topics such as continual learning and emergent intelligence. For Datasets, we cover existing benchmarks proposed in the AI community. They help bridge the gap between interdisciplinary theoretical definitions and the design of better evaluation benchmarks to pinpoint model capabilities. For Compositional Learning Models, we overview the architectures that aim to address compositional learning, divided into categories of basic neural architectures, large language models, and customized architectures, including neurosymbolic models. These models are mostly evaluated empirically using conventional benchmarks, while fewer studies conduct theoretical analyses. We cover both types of evaluations of various models when available in the existing literature. For Evaluation, we overview two evaluation approaches- theoretical and empirical. The theoretical evaluations examine various computational models in a mathematical framework, investigating their expressivity, and capacity for compositional learning, and analyze the generalizations to unobserved situations by computing the error bounds. Empirical evaluations include experimental results on benchmarks set by datasets and tasks created primarily to highlight the core challenges of compositional learning for language and vision understanding. Such results often report performance measures on the tasks designed to test the cognitive aspects of compositionality.
Overall, the cognitive aspects lay the foundation of the concept of compositionality and define the different abstractions associated with it and the tasks that are designed accordingly. The empirical evaluations use these tasks to evaluate compositionality using experimental performance. However, some studies examine the mathematical analysis and functional form of the models independent from the datasets. Finally, both these evaluations are used to develop models that are capable of compositional learning.
Figure 1: Outline of covered concepts in this survey, related to the structure of the paper. We structure our study of compositional learning by dividing it into four parts of compositional learning facets, datasets, compositional learning models, and evaluation methods from both empirical and theoretical perspectives. The topics have the respective sections associated with them. The main areas of required research and future direction are included in the descriptions in the Evaluation boxes, which are further discussed in Section 6.
@@ 2 Compositional Learning Facets
Compositionality is one important aspect of generalization as a whole (Hupkes et al., 2023). Cognitive Science and Linguistics literature have identified broad categorizations of tasks that define compositionality and can be used to evaluate the compositional reasoning of models. The foundations of human natural language lie in compositionality. A commonly used task categorization, derived from reformulated theoretically grounded tests from Hupkes et al. (2020), defines five main metrics of compositionality: systematicity, productivity, substitutivity, localism, and overgeneralization (Dankers et al., 2021).
2.1 Measures of Compositionality
Hupkes et al. (2020) introduces theoretically grounded tests for compositionality of models based on different interpretations of compositionality (Fodor & Pylyshyn, 1988; Chomsky, 1956). Some of these tasks existed in earlier research under different terms, such as productivity, which builds on a substantial body of prior research on length generalization. These tasks are becoming widely accepted tasks for compositional learning, and there are current datasets that use them for their evaluation splits. We describe these measures in detail below.
Systematicity or Novel Composition
Systematicity is one of the most commonly used notions of compositionality in evaluating the performance of computational models. It has been defined as the ability to systematically recombine known parts and rules (Dankers et al., 2021). It derives directly from the commonly accepted definition of composition, which is the formation of compound expressions as a function of simpler ones (Partee, 2004). Systematicity is a standard concept in cognitive science research on building cognitive architectures that try to formalize the human thinking process (Fodor & Pylyshyn, 1988). The ability to syntactically combine known elements to form new or “unseen” expressions is an integral test when evaluating a model’s ability to reason compositionally. It is also called Distribution-Based Compositionality Assessment (Keysers et al., 2020), where two principles are defined: one is to make sure the distribution of atoms is similar in both training and test sets, while another is to ensure the distribution of compounds is different. For example, if “red” and “car” are two separately learned concepts, the model should be able to accurately utilize the unseen concept of “red car.”
Productivity or Length Generalization
Another commonly explored test for compositionality is length generalization or productivity, as defined in Hupkes et al. (2020). In this evaluation, models are tested on their performance with expressions or sequences that are longer than training data. Longer input sequences may be recursive or nested versions of seen phrases in the case of natural language inputs (Kim & Linzen, 2020). For example, if the model has seen “A and B,” it should be able to understand “A and B and C.”
Substitutivity or Synonymity
Substitutivity is another form of evaluation defined in Dankers et al. (2021), which concerns the model performance on the introduction of synonyms in expressions. For example, testing a model on the translation of the same sentence, when switching between synonymous words such as donut to doughnut or aubergine to eggplant the translation would not change. This is one of the less explored axioms of compositionality.
Localism
Another nuance of compositionality is the notion of global versus local composition. According to the principle of compositionality, the locality of the composition operator can vary. The meaning of a complex expression can depend solely on the meaning of its immediate parts (local composition) or the global structure of the context. Localism can be tested by analyzing the meaning a model assigns to a standalone compound versus when that compound is part of a larger expression. For example, sentences X and Y have the same truth values, but if the same context is added, their composition with the new context might lead to different truth values. For instance, when we add the context, “Peter thinks” and obtain “Peter thinks X” and “Peter thinks Y,” these two new sentences can have different truth values (Hupkes et al., 2020; Carnap, 1947). The local interpretation of compositionality says that these new phrases will have the same truth values which might not be the case anymore as Peter might be aware of X and not Y. In other words, considering the phrase that X and Y are a part of, changes the meaning. This is another one of the less explored axioms of compositionality.
Overgeneralization
Overgeneralization, as defined in Dankers et al. (2021), evaluates how much a model prefers an exception versus a rule. The term is originally used in language acquisition literature, also known as overregularization. A well-known example of this is the past-tense debate (Marcus et al., 1992) in language, which is about the rule that English past-tense verbs can be formed by appending -ed to the stem of the verb in most cases while there are some exceptions. This property can be evaluated by testing a model on exceptions of a usual rule in the training data and seeing if the model has over-fitted the training samples (Hupkes et al., 2020). Another example of this task is translating idioms where the meaning of sentences is “exceptions” to usual rules. For example, when translating the idiom “it’s raining cats and dogs”, a literal translation does not make sense as the phrase is an exception and has a specific meaning different from the usual literal interpretation. In this scenario, a model can achieve better performance by considering sentences in a global sense -that is looking at the bigger picture, such as context from placement in a compound- instead of trying to evaluate the meaning locally, such as by isolating the phrase. This is yet another less evaluated axiom of compositionality.
2.2 Compositionality as Function Properties
In Ram et al. (2024), compositional functions are defined with several components with a computation directed acyclic graph (cDAG) at the core. Their formal definitions facilitate the evaluation of compositional properties of the learning models (i.e. compositional functions). They relate their defined structure to the learning models’ expressivity and sample complexity. In this system, Systematicity can be thought of as the expressivity of a compositional function as a low entropy program (for example, decision tree) versus a high entropy program (for example, transformer). Productivity, in simple terms, can be interpreted as whether a compositional function is recursive. Substitutivity tests whether a compositional function respects important abstractions and can be factored over them. Localism measures the stability of a compositional function against local changes, where the structure of the function’s elements affects the importance of the level of locality respected. Overgeneralization is the extent of compression of a function, where a function might have a general rule but have exceptions to those in special cases.
2.3 Compositionality and Continual Learning
Compositionality is an important aspect of continual learning, also known as lifelong learning (Mendez & Eaton, 2021). Continual learning (Wang et al., 2024) is the concept of learning new tasks while retaining knowledge from previously learned tasks. In continual learning, compositionality is particularly crucial to prevent catastrophic forgetting, where earlier tasks are forgotten over time due to learning of new tasks (Liao et al., 2023). Since knowledge about novel tasks is compositional, the existing information can be combined in novel ways and used for future tasks. This enables forward transfer of knowledge rather than catastrophic forgetting (Mendez & Eaton, 2023). Both compositional and continual learning share the key challenge of finding reusable knowledge and further connecting those for transfer learning and dealing with complex unobserved situations.
2.4 Compositionality and Emergent Intelligence
The term emergent behavior has been used across various science-related fields, rooted in “More Is Different” by Nobel Prize-winning physicist P.W. Anderson (Anderson, 1972). Its introduction to the language modeling community, specifically in the context of the large language models, begins with Wei et al. (2022). They defined emergent ability as the ability that appears only in large models and is not observed in any smaller models. The emergent abilities demonstrated in their study include performing unseen tasks by following instructions (Ouyang et al., 2024) and demonstrating multi-hop reasoning skills through Chain-of-Thought prompting (Wei et al., 2024). These abilities reflect the model’s capacity for generalizing to unobserved situations, which can further extend to the model’s compositional learning ability. Therefore, the compositional learning ability of the models can be associated with the emergent intelligence of language models. In the same paper of Wei et al. (2022), creating new compositional learning benchmarks is proposed as one direction for evaluating and understanding LLMs’ emergent abilities.
3 Abstract Tasks and Datasets
In this section, we categorize the existing datasets proposed for the evaluation of compositional learning. Our categorization is based on the type of compositionality facet explained in Section 2. Table 1 points to a list of important datasets we have surveyed. In general, there are more common evaluation benchmarks for Systematicity and Productivity. Systematicity focuses on the novel composition of seen atomic concepts and there are several benchmarks established for its evaluation, although some of those works do not explicitly use the term systematicity. The productivity measure was often referred to as length generalization before the term became commonly used. However, despite the abundance of datasets for these tasks, most rely on synthetic data, which poses a risk to their reliability in capturing the variation and complexity of real-world problems. While some datasets in the computer vision community for compositional learning utilize realistic images, they address fewer aspects of compositionality (e.g. object-attribute combination) compared to synthesized linguistic corpora designed for the same purpose. We describe some of the existing datasets and tasks in this section.
@@ 3.1 Systematicity or Novel Composition
CREPE. This is a Compositional REPresentation Evaluation benchmark (CREPE) (Ma et al., 2022). The dataset is synthesized and includes multiple splits, one of which relates to Systematicity. The main task setting is that, given an image, the model needs to identify an appropriate text caption describing it among multiple given choices. This systematicity challenge tests whether the model can systematically generate new combinations of seen atomic concepts during training. For example, “Crepe on a skillet” is never observed in the training while Crepe and skillet are observed separately in different contexts.
@@
SCAN. The task is to navigate in a two-dimensional grid world based on natural language instructions. It is the Simplified version of the CommAI Navigation tasks (SCAN) (Lake & Baroni, 2017; Mikolov et al., 2018). One of the proposed experiments in SCAN evaluates the model’s compositional generalization across primitive commands. Specific compounds are excluded from training where the model has seen the primitives and similar compound structures. Then, these unseen compounds are tested during testing.
@@
gSCAN. The task is to navigate in a two-dimensional grid world based on natural language instructions (Ruis et al., 2020), which is a grounded version of SCAN (gSCAN). It includes 9 splits A-I. Categories B to H present tasks that form tests for systematicity (B,C- Novel Composition of Object Properties, D- Novel Direction, E- Novel Contextual References, F- Novel composition of actions and arguments, G,H- Novel Adverbs). Each split focuses on some form of novel composition of known concepts.
@@
PCFG SET. PCFG SET stands for Probabilistic Context Free Grammar String Edit Task. It is an artificial translation task where sequences produced by probabilistic context-free grammar need to be translated into sequences representing their meaning (Hupkes et al., 2020). The output sequences can be constructed recursively using specified string edit operations applied to the input sequence, e.g., the input “repeat ABC” will be mapped to the sequence “ABCABC.” The systematicity test uses a combination of words $a$ and $b$ in the input where the model has seen $a$ but not $b$ and vice versa. However, the combination of $a$ and $b$ is plausible in the corpus.
@@
COGS. This dataset is designed for the Compositional Generalization Challenge (COGS) (Kim & Linzen, 2020). The task is a kind of semantic parsing based on a fragment of English where the models need to determine a formal meaning representation of the input English text. There are 4 categories in COGS that test some form of systematicity, including Novel Combination of Familiar Primitives and Grammatical Roles, Novel Combination of Modified Phrases and Grammatical Roles, Verb Argument Structure Alteration, and Verb Class.
ReaSCAN. This extension of gSCAN overcomes some of its limitations by requiring compositional language interpretation and reasoning about entities and relations (Wu et al., 2021). The challenges for systematicity are Category A, which tests novel object attribute combinations such as novel color modifier, color attribute, and size attribute, which is adapted from gSCAN (Ruis et al., 2020). Category B tests unseen co-occurrences of objects and relations, which is unique to ReaSCAN.
SQOOP. This is a dataset of Spatial Queries On Object Pairs (SQOOP). It is a minimalistic visual question answering, with yes-no answers. The input is an image and a question based on spatial reasoning (Bahdanau et al., 2019). Models are tested for answering questions on all possible object pairs after being trained on only a subset. Questions are of the form X R Y (X and Y are objects while R is a relation) where training sets are generated by controlled sampling of X and Y objects.
CLUTRR. This dataset is on Compositional Language Understanding and Text-based Relational Reasoning (CLUTRR). The task is, given a natural language short story, to answer questions on kinship relations that can be inferred from story (Sinha et al., 2019). Models are tested on combinations of held-out reasoning rules that are unseen during training. Thus, it tests systematic generalization capability or systematicity.
KiloGram. The task is a reference game task where given a textual description, the model has to select the appropriate image from a set of images. These images are tangrams, and the dataset has rich annotations of these images in two splits, FULL and DENSE, which have varying numbers of annotations, hence the name Kilo Tangram KiloGram) (Ji et al., 2022). There are different variations of this task such as showing parts of the image versus the whole image and making it grayscale versus colored. This dataset is an example of compositionality in vision since the whole tangram image is made out of parts, and the model learns how different parts combine to form different images.
CompMCTG. To evaluate the compositional learning of generative language models, Compositional multi-aspect controllable text generation (CompMCTG) benchmark(Zhong et al., 2024) is proposed. The task is to generate sentences given a set of concepts, including aspects of sentiment, topic, tense, person, and stuff. The benchmark tests systematicity by evaluating the model’s capability to generate sentences with novel or unseen combinations of such concepts. For example, if the model has seen sentences with concepts of (red, car) and (blue, hat), it should be able to generate sentences for (blue, car). To perform these evaluations, they split the dataset into in-distribution (I.D.) and compositional, which have no intersection, although recombining elements in the I.D. set can form elements in the compositional set. Their evaluation protocol includes three test splits of Hold-Out, Attribute Compound Divergence (ACD), and Few-Shot evaluation. It is based on four popular textual corpora- Amazon Reviews (He & McAuley, 2016), a combination of IMDB, OpeNER, SenTube: Mixture (Liu et al., 2022a), YELP (Yelp, 2014), and Fyelp (Lample et al., 2019).
MIT-States. In this evaluation benchmark (Isola et al., 2015), the general task is to explain a collection of images in terms of the novel composition of primitive states and transformations applied to objects. It includes three different tasks: discovering relevant transformations (such as slicing, wilting), parsing states (such as sliced, wilted), and finding smooth transitions. It tests systematicity in image concepts such as being able to identify “sliced tomato” versus “whole tomato” and generalize it to unobserved compositions such as “sliced apple” versus “whole apple.” There exists other similar works (Yu & Grauman, 2014; 2017; Tokmakov et al., 2019; Xu et al., 2023b; Bao et al., 2024; Xu et al., 2023a) in the vision community that evaluate such compositional concept learning.
Skill-Mix. There is no commonly used benchmark for the evaluation of compositional learning abilities of LLMs. A very recent benchmark, called Skill-Mix (Yu et al., 2024) covers some aspects of compositional evaluation for generative models. The task expects the models to generate text by combining various skills and imposing some constraints on the generated text. While the notion of compositionality is not highlighted in the paper, we categorize their tasks under systematicity.